Schedule Y-1-Married Filing Jointly or Qualifying Widow(er) If taxable income is over: But not over: The tax is: $0 $19,400 $Determine the amount of tax liability in the following situations. In all cases, the taxpayer is using the filing status of married filing jointly. Use the appropriate Tax Tables or Tax Rate Schedules.

  1. Taxable income of $62,449 that includes a qualified dividend of $560.
  2. Taxable income of $12,932 that includes a qualified dividend of $322.
  3. Taxable income of $144,290 that includes a qualified dividend of $4,384. (Round your intermediate computations to 2 decimal places and final answer to the nearest whole dollar amount.)
  4. Taxable income of $43,297 that includes a qualified dividend of $971.
  5. Taxable income of $262,403 that includes a qualified dividend of $12,396. (Round your intermediate computations to 2 decimal places and final answer to the nearest whole dollar amount.)
Answer #1
  • Taxable Income = $(62,449-560) = $61,889

Tax liability will be $1,940 + 12% of the excess over $19,400

Excess over $19,400 = $(61,889-19,400) = $42,489

Therefore, Tax Liability = $1,940 + $42,489*12% = $(1,940+5,098.68) = $7,039 (rounded off)

  • Taxable Income = $(12,932-322) = $12,610

Tax liability will be 10% of taxable income = $12,610*10% = $1,261

  • Taxable Income = $(144,290-4,384) = $139,906

Tax liability will be $9,086 + 22% of the excess over $78,950 + 15% of qualified dividend

Excess over $78,950 = $(139,906-78,950) = $60,956

Therefore, Tax Liability = $9,086 + $60,956*22% + $4,384*15% = $(9,086+13,410.32+657.6) = $23,154 (rounded off)

  • Taxable Income = $(43,297-971) = $42,326

Tax liability will be $1,940 + 12% of the excess over $19,400

Excess over $19,400 = $(42,326-19,400) = $22,926

Therefore, Tax Liability = $1,940 + $22,926*12% = $(1,940+2,751.12) = $4,691 (rounded off)

  • Taxable Income = $(262,403-12,396) = $250,007

Tax liability will be $28,765 + 24% of the excess over $168,400 + 15% of qualified dividend

Excess over $168,400 = $(250,007-168,400) = $81,607

Therefore, Tax Liability = $28,765 + $81,607*24% + $12,396*15% = $(28,765+19,585.68+1,859.4) = $50,210 (rounded off)